OpenStudy (anonymous):

what's the derivative of the square root of ((x^2)*(6^x))^9?

OpenStudy (anonymous):

\[\frac{d}{dx}(\sqrt{x^26^x})^9\] like that?

OpenStudy (lukecrayonz):

0

OpenStudy (anonymous):

the 9 is under the square root

OpenStudy (anonymous):

to find the derivative of that.

OpenStudy (anonymous):

\[\sqrt{((x^2)*(6^x))^9}\] sorry it's actually this

OpenStudy (anonymous):

hey u have to use the rule of partial differentiatio i.e., (d/dx)(x^18*6^9x)

OpenStudy (anonymous):

maybe best to write \[x^2x^6=x^8\] and \[(x^8)^9=x^{72}\] and \[\sqrt{x^{72}}=x^{36}\] and take the derivative of that

OpenStudy (anonymous):

but because it's 6^x and not x^6 i don't think i can do that

OpenStudy (anonymous):

ooooh ok then you have to do something else

OpenStudy (anonymous):

\[\sqrt{(x^2\times 6^x)^9}\] like this?

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{x^{18}6^{9x}}=\frac{1}{2\sqrt{x^{18}6^{9x}}}\times \frac{d}{dx}[x^{18}6^{9x}]\] by the chain rule

OpenStudy (anonymous):

then \[\frac{d}{dx}[x^{18}6^{9x}]=18x^{17}6^{9x}+x^{18}\times 9\times \ln(x)\times 6^{9x}\]

OpenStudy (anonymous):

then some algebra to clean it up